OpenStudy (anonymous):

Compute d2y/dx2 at the point(2,2) x^2+y^2=8

OpenStudy (anonymous):

Can you take dy/dx?

OpenStudy (anonymous):

i did take dy/dx. i got 2x+ 2y dy/dx=0 that led to -2x/2y=-1

OpenStudy (anonymous):

i dont know how to do it the 2nd time

OpenStudy (anonymous):

It's correct dy/dx = - x/y

OpenStudy (anonymous):

hmm k i get -y + xdy/dx/y^2

OpenStudy (anonymous):

(y + xdy/dx)/y^2

OpenStudy (anonymous):

wait is the answer 3

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

but its not any of the answer choices

OpenStudy (anonymous):

the choices are a)-3 b)1 c)2 d) -1 e)-2

OpenStudy (anonymous):

Hold, I think you have wrong sign!

OpenStudy (anonymous):

= 2 - 2 ( -1) / 2^2 = 4/4 = 1

OpenStudy (anonymous):

hmm k. i hope its correct

OpenStudy (anonymous):

Math has the numbers, it's NOT hope!

OpenStudy (anonymous):

You're correct at derivative, but somehow you miscalculated at last step!

OpenStudy (anonymous):

k. thanks!

OpenStudy (anonymous):

it was wrong :(

OpenStudy (anonymous):

y'=(-x/y) then y''=(-y+xy')/y^2 then y''=(-y-x^2/y)/y^2=-(x^2+y^2)/y^3=-8/y^3=-1

OpenStudy (anonymous):

@mahmit2012's result is right! I have worked out on it to check the plug in : y" = ( -y + xy' ) / y^2 = -2 +2 ( -1) / 4 = - 4/ 4 = -1

OpenStudy (anonymous):

yes but i wanted to show in this case (circle equation)we have y''=-R^2/y^3 R is radiant of circle